OpenStudy (anonymous):

Sarah is driving her car very smoothly from Austin to Dallas one day. At 6:00 pm she is travelling at 20 mph, while 25 minutes later she is travelling at 40 mph. Calculus tells us that at some time between 6:00 pm and and 6:25 pm her acceleration is exactly x mph/h. Determine x.

OpenStudy (anonymous):

this is the mean value theorem

OpenStudy (anonymous):

like saying \[v(0)=20, v(25)=40\] and so by mean value theorem there is an x between 0 and 25 with \[v'(x)=\frac{40-20}{25}=\frac{20}{25}=\frac{4}{5}\]]

OpenStudy (anonymous):

and since the derivative of the velocity is the acceleration, that is your answer

OpenStudy (anonymous):

Yeah. What satellite says. 4/5 mph/min but we want x in mph/h, so multiply by 60 (4/5)*60=48 mph/h

OpenStudy (anonymous):

Thanks both ya

OpenStudy (anonymous):

oh good point!